# Planar Bilipschitz Extension from Separated Nets

### v2

@@ Acknowledgements.
Let $X\subseteq\mathbb{R}^{2}$. For each $x\in X$ let $y_{x}\in\mathbb{R}^{2}$, $r_{x}$ be a positive real number and $\mathfrak{U}_{x}:=B([x,y_{x}],r_{x})$. Suppose that $(\mathfrak{U}_{x})_{x\in X}$ is pairwise disjoint,
$\sup_{x\in X,\,y_{x}\neq x}\frac{r_{x}}{\left\|y_{x}-x\right\|}\leq\frac{1}{2}\qquad\text{and}\qquad\sup_{x\in X,\,y_{x}\neq x}\frac{\left\|y_{x}-x\right\|^{2}}{r_{x}^{2}}<\infty.$
@@
We may now establish our main result, Theorem 1.1, by verifying for the given bilipschitz $f\colon\mathbb{Z}^{2}\to\mathbb{R}^{2}$ the condition (ii) of Theorem 1.2. This task can be broken down into two steps: First extend $f$ on a sequence of separated horizontal lines partitioning $\mathbb{R}^{2}$ into large strips. That is, extend $f$ to $\mathbb{Z}^{2}\cup\left(\mathbb{R}\times\left(T\mathbb{Z}+\frac{1}{2}\right)\right)$ for $T\in\mathbb{N}$ chosen suitably large, to obtain the mapping $\widetilde{F}$. Next, restrict $\widetilde{F}$ to the horizontal lines $\mathbb{R}\times\left(T\mathbb{Z}+\frac{1}{2}\right)$, ‘throwing away $\mathbb{Z}^{2}$ and $f$’, and extend this mapping to the whole of $\mathbb{R}^{2}$ to get $G$. The first step in this strategy, the extension of $f$ to $\widetilde{F}$, is the most difficult and requires preparation in all subsequent sections of the paper, before it is completed in Section 5: {restatable*}thmrestatestrips Let $L\geq 1$, $T\in\mathbb{N}$, $T\geq 10^{38}L^{16}$ and $f\colon\mathbb{Z}^{2}\to\mathbb{R}^{2}$ be an $L$-bilipschitz mapping. Then there exists a $10^{37}L^{15}$-bilipschitz extension
@@
of $f$ such that $\operatorname{bilip}\left(F|_{\mathbb{R}\times\left(T\mathbb{Z}+\frac{1}{2}\right)}\right)\leq 10^{23}L^{11}$ and, letting $V_{i}$, for each $i\in\mathbb{Z}$, denote the (unique) connected open subset of $\mathbb{R}^{2}$ with boundary equal to $F\left(\mathbb{R}\times\left\{T(i-1)+\frac{1}{2},Ti+\frac{1}{2}\right\}\right)$, the following statements hold:
@@
$T:=\lceil 10^{38}L^{16}\rceil,\qquad M_{1}:=10^{37}L^{15},\qquad K:=10^{23}L^{11},\quad\text{and the}\\ M_{1}\text{-bilipschitz mapping }\widetilde{F}\colon\mathbb{Z}^{2}\cup\left(\mathbb{R}\times\left(T\mathbb{Z}+\frac{1}{2}\right)\right)\to\mathbb{R}^{2}\text{ with }\operatorname{bilip}\left(\widetilde{F}|_{\mathbb{R}\times\left(T\mathbb{Z}+\frac{1}{2}\right)}\right)\leq K,$
@@
where $V_{i}$ for $i\in\mathbb{Z}$ is defined in the statement of Theorem 1.3. The proof is now completed by Theorem 1.2: observe that $T$, $G$ verify statement (ii) of Theorem 1.2 and $M_{1}$, $M_{2}$, the hypothesis of its ‘Furthermore’ statement. To refine the upper bound on $\operatorname{bilip}(F)$ from Theorem 1.2 to the form in the present statement, we apply $T\leq 10^{40}L^{16}$. ∎
@@
We write :=:= to signify a definition by equality. For $n\in\mathbb{N}$ the notation $[n]$ refers to $\left\{1,\ldots,n\right\}$. Moreover, for $P\in\mathbb{R}$, the notation $\mathbb{Z}_{\geq P}$ (e.g.) stands for $\left\{i\in\mathbb{Z}\colon i\geq P\right\}$. We write $e_{1},e_{2}$ for the standard basis of $\mathbb{R}^{2}$. The orthogonal projection to the $i$-th coordinate in $\mathbb{R}^{2}$ is denoted by $\operatorname{proj}_{i}$ for $i=1,2$.
@@
Given $A\subset\mathbb{R}^{2}$ and $r>0$, we say that $A$ is $r$-separated if $\left\|a-a^{\prime}\right\|\geq r$ for every $a^{\prime}\in A,a\neq a^{\prime}$. We say that $A$ is an $r$-net if $d(x,A)\leq r$ for every $x\in\mathbb{R}^{2}$. If there are $r,s>0$ such that $A$ is $s$-separated $r$-net, we call $A$ a separated net.
@@
We set $f:=f_{1}\cup f_{2}$. The assumptions ensure that $f^{-1}$ is well-defined and that $A_{1}$ is closed. Since $f_{1}$ is injective, by Brouwer’s Invariance of Domain [8, Thm 2B.3] it holds that $f_{1}(\partial A_{1})=\partial f_{1}(A_{1})$. Consequently, $\partial f_{1}(A_{1})\subseteq f_{1}(A_{1}\cap A_{2})=f_{1}(A_{1})\cap f_{2}(A_{2})$. Thus, the assumptions are symmetric with respect to $(f_{1},f_{2},f)$ and $(f_{1}^{-1},f_{2}^{-1},f^{-1})$ and it suffices to show that $f$ is $L$-Lipschitz, as follows: Given any $x_{1}\in A_{1},x_{2}\in A_{2}\setminus A_{1}$ we take $z(x_{1},x_{2})$ as any point of $x_{1},x_{2}]\cap\partial A_{1}$ and invoke Lemma 4 with $C=1$. ∎
@@
Let $L\geq 1$, $k\in\mathbb{Z}$ and $f\colon\mathbb{Z}^{2}\to\mathbb{R}^{2}$ be an $L$-bilipschitz mapping. Then there exists a $10^{38}L^{15}$-bilipschitz extension
@@
Let $L$, $T$ and $f$ be given by the hypotheses of Theorem 1.3 and let $F_{i}$, for each $i\in\mathbb{Z}$, be given by the conclusion of Lemma 7 with $k=Ti$. Then $F:=\bigcup_{i\in\mathbb{Z}}F_{i}\colon\mathbb{Z}^{2}\cup\left(\mathbb{R}\times\left(T\mathbb{Z}+\frac{1}{2}\right)\right)\to\mathbb{R}^{2}$ is well-defined and $T$ is large enough to ensure $F$ is $10^{38}L^{15}$-bilipschitz with $\operatorname{bilip}\left(F|_{\mathbb{R}\times\left(T\mathbb{Z}+\frac{1}{2}\right)}\right)\leq 10^{24}L^{11}$; to verify these constants apply [6, Lemma 6.2], with $d=2$, $J=10^{38}L^{15}$ and $K=10^{24}L^{11}$. Finally (i) and (ii) follow from the definition of $F$ and the last stated property of $F_{i}$ from the conclusion of Lemma 7; this is confirmed by [6, Lemma 6.3]. ∎
@@
and let $r_{x}:=\frac{\eta}{5}$ and $\mathfrak{U}_{x}:=B([x,y_{x}],r_{x})$. Since $\Phi\circ f(\mathbb{Z}^{2})$ is $\eta$-separated, and $\mathfrak{U}_{x}\subseteq B(x,\left\|y_{x}-x\right\|+r_{x})\subseteq B(x,\eta/2)$ for each $x\in\Phi\circ f(\mathbb{Z}^{2})$, we have that the family $(\mathfrak{U}_{x})_{x\in\Phi\circ f(\mathbb{Z}^{2})}$ is pairwise disjoint. Let the $4$-bilipschitz mappping $\Gamma\colon\mathbb{R}^{2}\to\mathbb{R}^{2}$, with
@@
$\Upsilon\circ\xi(t)=(t,0)\quad\text{for all }t\in\mathbb{R},\quad\text{ and}\qquad\operatorname{bilip}(\Upsilon)\leq 4\cdot 10^{5}L.$
@@
$\operatorname{bilip}(\Upsilon^{-1}\circ\gamma)$ $\leq 4\cdot 10^{5}L\cdot\frac{1}{4}10^{18}L^{10}=10^{23}L^{11}=:M,$
@@
$\geq\frac{1}{4\cdot 10^{5}L}\cdot\frac{4}{10^{9}L^{2}}=\frac{1}{10^{14}L^{3}}=:\theta\quad\text{for all $x\in\mathbb{Z}^{2}$,}\quad\text{and}$
@@
$\leq 4\cdot 10^{5}L\cdot\left(1+2\cdot 10^{7}L^{8}\right)+8L^{5}\leq 2\cdot 10^{13}L^{9}=:K\quad\text{for all $i\in\mathbb{Z}$,}$
allow us to apply Lemma 9 to establish $\operatorname{bilip}(F)\leq\frac{ML}{\theta}=10^{37}L^{15}$. We also get $\operatorname{bilip}\left(F|_{\mathbb{R}\times\left\{\frac{1}{2}\right\}}\right)=\operatorname{bilip}(\Upsilon^{-1}\circ\gamma)\leq 10^{23}L^{11}$. ∎

### v3

@@ Acknowledgements.
Let $X\subseteq\mathbb{R}^{2}$. For each $x\in X$ let $y_{x}\in\mathbb{R}^{2}$, $r_{x}$ be a positive real number and $\mathfrak{U}_{x}:=B([x,y_{x}],r_{x})$ (the set of points in $\mathbb{R}^{2}$ with distance less than $r_{x}$ to the line segment $x,y_{x}]$; see Section 2). Suppose that $(\mathfrak{U}_{x})_{x\in X}$ is pairwise disjoint,
$\sup_{x\in X,\,y_{x}\neq x}\frac{r_{x}}{\left\|y_{x}-x\right\|}\leq\frac{1}{2}\qquad\text{and}\qquad\sup_{x\in X,\,y_{x}\neq x}\frac{\left\|y_{x}-x\right\|}{r_{x}}<\infty.$
@@
We may now establish our main result, Theorem 1.1, by verifying for the given bilipschitz $f\colon\mathbb{Z}^{2}\to\mathbb{R}^{2}$ the condition (ii) of Theorem 1.2. This task can be broken down into two steps: First extend $f$ on a sequence of separated horizontal lines partitioning $\mathbb{R}^{2}$ into large strips. That is, extend $f$ to $\mathbb{Z}^{2}\cup\left(\mathbb{R}\times\left(T\mathbb{Z}+\frac{1}{2}\right)\right)$ for $T\in\mathbb{N}$ chosen suitably large, to obtain the mapping $\widetilde{F}$. Next, restrict $\widetilde{F}$ to the horizontal lines $\mathbb{R}\times\left(T\mathbb{Z}+\frac{1}{2}\right)$, ‘throwing away $\mathbb{Z}^{2}$ and $f$’, and extend this mapping to the whole of $\mathbb{R}^{2}$ to get $G$. The first step in this strategy, the extension of $f$ to $\widetilde{F}$, is the most difficult and requires preparation in all subsequent sections of the paper, before it is completed in Section 5: {restatable*}thmrestatestrips Let $L\geq 1$, $T\in\mathbb{N}$, $T\geq 10^{40}L^{16}$ and $f\colon\mathbb{Z}^{2}\to\mathbb{R}^{2}$ be an $L$-bilipschitz mapping. Then there exists a $10^{39}L^{15}$-bilipschitz extension
@@
of $f$ such that $\operatorname{bilip}\left(F|_{\mathbb{R}\times\left(T\mathbb{Z}+\frac{1}{2}\right)}\right)\leq 10^{24}L^{11}$ and, letting $V_{i}$, for each $i\in\mathbb{Z}$, denote the (unique) connected open subset of $\mathbb{R}^{2}$ with boundary equal to $F\left(\mathbb{R}\times\left\{T(i-1)+\frac{1}{2},Ti+\frac{1}{2}\right\}\right)$, the following statements hold:
@@
$T:=\lceil 10^{40}L^{16}\rceil,\qquad M_{1}:=10^{39}L^{15},\qquad K:=10^{24}L^{11},\quad\text{and the}\\ M_{1}\text{-bilipschitz mapping }\widetilde{F}\colon\mathbb{Z}^{2}\cup\left(\mathbb{R}\times\left(T\mathbb{Z}+\frac{1}{2}\right)\right)\to\mathbb{R}^{2}\text{ with }\operatorname{bilip}\left(\widetilde{F}|_{\mathbb{R}\times\left(T\mathbb{Z}+\frac{1}{2}\right)}\right)\leq K,$
@@
where $V_{i}$ for $i\in\mathbb{Z}$ is defined in the statement of Theorem 1.3. The proof is now completed by Theorem 1.2: observe that $T$, $G$ verify statement (ii) of Theorem 1.2 and $M_{1}$, $M_{2}$, the hypothesis of its ‘Furthermore’ statement. To refine the upper bound on $\operatorname{bilip}(F)$ from Theorem 1.2 to the form in the present statement, we apply $T\leq 10^{41}L^{16}$. ∎
@@
We use largely a common notation in the present article and its companion article [6]. We write :=:= to signify a definition by equality. For $n\in\mathbb{N}$ the notation $[n]$ refers to $\left\{1,\ldots,n\right\}$. Moreover, for $P\in\mathbb{R}$, the notation $\mathbb{Z}_{\geq P}$ (e.g.) stands for $\left\{i\in\mathbb{Z}\colon i\geq P\right\}$. We write $e_{1},e_{2}$ for the standard basis of $\mathbb{R}^{2}$. The orthogonal projection to the $i$-th coordinate in $\mathbb{R}^{2}$ is denoted by $\operatorname{proj}_{i}$ for $i=1,2$.
@@
Given $A\subset\mathbb{R}^{2}$ and $r>0$, we say that $A$ is $r$-separated if $\left\|a-a^{\prime}\right\|\geq r$ for every $a,a^{\prime}\in A,a\neq a^{\prime}$. We say that $A$ is an $r$-net if $\operatorname{dist}(x,A)\leq r$ for every $x\in\mathbb{R}^{2}$. If there are $r,s>0$ such that $A$ is an $s$-separated $r$-net, we call $A$ a separated net.
@@
We set $f:=f_{1}\cup f_{2}$. The assumptions ensure that $f^{-1}$ is well-defined and that $A_{1}$ is closed. Since $f_{1}$ is bilipschitz, $f_{1}(A_{1})$ is also closed, so $\partial f_{1}(A_{1})\subseteq f_{1}(A_{1})$. Using Brouwer’s Invariance of Domain [8, Thm 2B.3], we conclude that $\partial f_{1}(A_{1})\subseteq f_{1}(\partial A_{1})$. Hence $\partial f_{1}(A_{1})\subseteq f_{1}(A_{1}\cap A_{2})=f_{1}(A_{1})\cap f_{2}(A_{2})$. Thus, the assumptions are symmetric with respect to $(f_{1},f_{2},f)$ and $(f_{1}^{-1},f_{2}^{-1},f^{-1})$ and it suffices to show that $f$ is $L$-Lipschitz, as follows: Given any $x_{1}\in A_{1},x_{2}\in A_{2}\setminus A_{1}$ we take $z(x_{1},x_{2})$ as any point of $x_{1},x_{2}]\cap\partial A_{1}$ and invoke Lemma 4 with $C=1$. ∎
@@
Let $L\geq 1$, $k\in\mathbb{Z}$ and $f\colon\mathbb{Z}^{2}\to\mathbb{R}^{2}$ be an $L$-bilipschitz mapping. Then there exists a $10^{39}L^{15}$-bilipschitz extension
@@
Let $L$, $T$ and $f$ be given by the hypotheses of Theorem 1.3 and let $F_{i}$, for each $i\in\mathbb{Z}$, be given by the conclusion of Lemma 7 with $k=Ti$. Then $F:=\bigcup_{i\in\mathbb{Z}}F_{i}\colon\mathbb{Z}^{2}\cup\left(\mathbb{R}\times\left(T\mathbb{Z}+\frac{1}{2}\right)\right)\to\mathbb{R}^{2}$ is well-defined and $T$ is large enough to ensure $F$ is $10^{39}L^{15}$-bilipschitz with $\operatorname{bilip}\left(F|_{\mathbb{R}\times\left(T\mathbb{Z}+\frac{1}{2}\right)}\right)\leq 10^{24}L^{11}$; to verify these constants apply [6, Lemma 6.2], with $d=2$, $J=10^{39}L^{15}$ and $K=10^{24}L^{11}$. Finally (i) and (ii) follow from the definition of $F$ and the last stated property of $F_{i}$ from the conclusion of Lemma 7; this is confirmed by [6, Lemma 6.3]. ∎
@@
and let $r_{x}:=\frac{\eta}{10}$ and $\mathfrak{U}_{x}:=B([x,y_{x}],r_{x})$. Since $\Phi\circ f(\mathbb{Z}^{2})$ is $\eta$-separated, and $\mathfrak{U}_{x}\subseteq B(x,\left\|y_{x}-x\right\|+r_{x})\subseteq B(x,\eta/2)$ for each $x\in\Phi\circ f(\mathbb{Z}^{2})$, we have that the family $(\mathfrak{U}_{x})_{x\in\Phi\circ f(\mathbb{Z}^{2})}$ is pairwise disjoint. Let the $16$-bilipschitz mapping $\Gamma\colon\mathbb{R}^{2}\to\mathbb{R}^{2}$, with
@@
$\Upsilon\circ\xi(t)=(t,0)\quad\text{for all }t\in\mathbb{R},\quad\text{ and}\qquad\operatorname{bilip}(\Upsilon)\leq 4^{2}\cdot 10^{5}L.$
@@
$\operatorname{bilip}(\Upsilon^{-1}\circ\gamma)$ $\leq 4^{2}\cdot 10^{5}L\cdot\frac{1}{4}10^{18}L^{10}\leq 10^{24}L^{11}=:M,$
@@
$\geq\frac{1}{4^{2}\cdot 10^{5}L}\cdot\frac{4}{10^{9}L^{2}}\geq\frac{1}{10^{15}L^{3}}=:\theta\quad\text{for all $x\in\mathbb{Z}^{2}$,}\quad\text{and}$
@@
$\leq 4^{2}\cdot 10^{5}L\cdot\left(1+2\cdot 10^{7}L^{8}\right)+8L^{5}\leq 10^{14}L^{9}=:K\quad\text{for all $i\in\mathbb{Z}$,}$
allow us to apply Lemma 9 to establish $\operatorname{bilip}(F)\leq\frac{ML}{\theta}=10^{39}L^{15}$. We also get $\operatorname{bilip}\left(F|_{\mathbb{R}\times\left\{\frac{1}{2}\right\}}\right)=\operatorname{bilip}(\Upsilon^{-1}\circ\gamma)\leq 10^{24}L^{11}$. ∎